A very cold but fine day at Grosmont today, with Ian Pearson, Steve Hyman, Nigel Bill, myself and later Jon Bradley working on the T2 (and Bill Dobson who called in early to brief us on what jobs to get on with) while Paul Hutchinson, Angie Buxton, and Mike Bloomfield were working on the K1 with Chris Henwood and Richard De Sadeleer joining them on the late shift.
No work was done on the J27, with the left hand valve repair being dealt with by the MPD staff and an instruction for the locomotive to be left alone for them to deal with. It will be in use however, for the Northern Light Express and the Santa Special trains from the end of November to the end of December, being based at Pickering to also provide steam heating for the duration. Barney Casey will be keeping an eye on it while it is there, and making sure it is not affected by any bad weather that might occur.
First job was to clean out the ashes from the stove, get it lit up and then have a warming cuppa. My first task was to clean out the remaining ash from the T2 smoke box so that Ian could get on with removal of the second main steam pipe. Then Steve, with my assistance, removed and cleaned all the cylinder cocks and stored them upstairs. The operating linkages were tied up and left in situ. Nigel Bill removed the valve that feeds steam to the reverser (this is situated on the right hand side of the boiler) after the application of heat to free up the securing nuts. The valve was dismantled in the workshop and then taken down to the MPD and given to Owain to do some machining and rectification, as it was blowing by when the carrot valve was open while topping up with steam oil.
Ian worked all day removing the nuts from the L/H main steam pipe in the smokebox with help from Nigel. This involved removal of a fair amount of concrete which was covering some of the nuts. By the end of the day, the steam pipe was successfully removed. Jon removed the injector caps for frost protection, cleaned them and put them upstairs. He also refitted all the cylinder head nuts back onto the studs after cleaning the threads and oiling them. The T2 is due a wash out, but this will be done as and when it becomes accessible for a move down to the MPD.
On the K1, there was frustratingly slow progress. The front valve was cleaned and fitted by Angie and Paul, the tail rod lubricator pipes annealed by Paul who also adjusted, fitted and secured the piston valve glands and made and fitted a new securing key for the R/H piston valve cotter. Both clack boxes were stripped for examination by Angie and she cleaned and lapped in both clack valves. Mike completed the cleaning of the R/H piston, and made a start on cleaning the R/H piston packing assembly. Late on in the day, Chris Henwood and Richard De Sadeleer made further progress with the replacement of the wasted roof stay nuts. In addition, Mark O’Brien made a start on re-profiling the throat plate mud hole door openings under contract with the NYMR. Paul also discussed with Barney the arranagements for getting the Support Coach down to Pickering at the end of the season, and also the use of the K1 during the Santa season out of Grosmont.
Working parties will continue tomorrow (15th) and Saturday (16th - when the JVs will also be present). On that Saturday, Mike McPeake is also running an MPD site briefing from 10.00 to 12.00 in the MIC Room at the MPD. This is essentially for the JVs, but if anyone else has not so far attended a briefing, or needs a refresher, and is able to go on Saturday they will be most welcome.
There will be a further working party on Wednesday 20 November, for which Neil Smedley and Ian McCall have already indicated their intention to be present.